1. A bullet moving at 200 m/sec strikes a tree and penetrates a distance of 0.75 cm into the tree before stopping. What was the deceleration of the bullet? How long did it take the bullet to stop? Assume constant deceleration.
Answer #1

Using second equation of motion,

a = v^2/2s = 200^2/[2*0.75e-2] = 2666667 m/s^2

time taken = v/a = 200/2666667 = 0.000075 s